# Localized frames without inequalities

### v1

@@ 1. Introduction
In many settings, however, frame theoretic results are ultimately not the consequence of an underlying group structure, but follow from localization properties of the frame, see, e.g., kernel theorems (compare [7, 15] with [11]), invertibility results of Gabor frame operators on modulation spaces [20, 24, 26], or a partial generalization [26, Theorem 5.1] of [22, Theorem 3.1]. A frame $(\psi_{k})_{k\in X}$ is said to be localized, if its Gram matrix is contained in a spectral matrix algebra (see [16] or Definition 2). To a localized frame $\psi$ one can then associate a family of Banach spaces, the so-called co-orbit spaces ${\mathcal{H}}^{p}(\psi)\ (1\leq p\leq\infty)$ which are defined in terms of summability conditions on the analysis coefficients $(\langle f,\psi_{k}\rangle)_{k\in X}$, see Section 2.3 for further details.
Returning to Gabor analysis, we note that the property $g\in M^{1}(\mathbb{R}^{d})$ of the window function can be seen as a localization assumption. Indeed, in the case $\Lambda=\mathbb{Z}^{d}$, we note that ${\mathcal{G}}(g,\mathbb{Z}^{d})$ is localized with respect to the unweighted Baskakov-Gohberg-Sjöstrand algebra $\mathcal{C}$ (see Example 4 (3)), and in the case of more irregular sets $\Lambda$, similar (more general) properties have been shown [26]. Furthermore, if ${\mathcal{G}}$ is a frame, the modulation spaces $M^{p}({\mathbb{R}}^{d})$ coincide with the co-orbit spaces $\mathcal{H}^{p}({\mathcal{G}})$ ($1\leq p\leq\infty$) generated from ${\mathcal{G}}$ in this case [26, Theorem 3.2]. Naturally, these observations lead to the question, whether abstract localized frames can be characterized in a similar fashion as in [22], namely via properties of their associated frame-related operators on their respective co-orbit spaces.
@@
Second, to our knowledge, there is currently no canonical generalization of the Ron-Shen duality principle to abstract frames in the literature. The Ron-Shen duality principle connects any Gabor frame on a lattice with its (dual) Gabor Riesz sequence generated by the same window on the adjoint lattice. In Theorem 1 we consider the family $\omega$, which, in the language of [33], is a so-called R-dual of type III of $\psi$, and serves as a substitute of the aforementioned Ron-Shen dual. In general, however, an R-dual associated with a Gabor frame on a lattice does not coincide with its Ron-Shen dual, see [12, 33] for a discussion of this matter. In order to guarantee that the R-dual $\omega$ is a Riesz sequence if and only if $\psi$ is a frame, we additionally have to assume that the reference system $\varphi$ is in fact a Riesz basis and not only a frame. Thus, our localization assumption $\psi\sim_{\mathcal{A}}\varphi$ corresponds to the original localization concept from [20] in the case ${\mathcal{A}}={\mathcal{J}}_{s}$ (see Example 4 (1)), which was followed by several other publications on this topic [1, 4, 5, 6, 16, 21]. At the same time, however, it might be very difficult, if not even impossible for a Gabor frame to be localized with respect to a localized Riesz basis (see also [20, Section 5] for a discussion of this issue).
Finally, several closed range conditions appear in our setting but not in [22, Theorem 3.1]. We show in Example 28 and Remark 29, that each of the closed range conditions appearing in Theorem 1 is in fact necessary in our setting, since otherwise, the statement is simply not true. For Gabor frames, however, these conditions are automatically satisfied, see Example 27.
By the previous observations, Theorem 1 does not directly generalize [22, Theorem 3.1]. However, [22] served as inspiration for this article, and multiple equivalent conditions of [22, Theorem 3.1] can be translated directly into conditions in Theorem 1. We refer to Section 5 for a more detailed comparison.
This article is structured as follows. In Section 2 we fix our notation and summarize some relevant concepts and results from the literature. In Section 3 we prove some auxiliary results, which we utilize in Section 4, where we prove our main result. In Section 5 we compare Theorem 1 to [22, Theorem 3.1], which initiated the research on the content of this article. Finally, Section 6 is devoted to applying our main result in the setting of shift-invariant spaces.
@@ 2.2. The R-dual sequence
converges in $\mathcal{H}$ for each $k\in X$ and the family $\omega=(\omega_{k})_{k\in X}$ is called an R-dual or Riesz-dual sequence for $\varphi$. It was shown in [12, Propositions 10,11] that $\varphi$ is a frame for ${\mathcal{H}}$ if and only if $\omega$ is a Riesz sequence.
@@ 2.3. Localized Frames
A unital Banach *-algebra $\mathcal{A}$ of $\mathbb{C}$-valued matrices $A=[A_{k,l}]_{k,l\in X}$ is called a solid spectral matrix algebra [16, Section 2.3], or simply spectral algebra, if
@@
Condition (A2) is equivalent to spectral invariance of ${\mathcal{A}}$ in ${\mathcal{B}}(\ell^{2}(X))$, that is, $\sigma_{{\mathcal{A}}}(A)=\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)$ for all $A\in{\mathcal{A}}$ [23, Lemma 2.4], where $\sigma_{{\mathcal{A}}}(A)$ and $\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)$ denote the spectrum of $A$ in the Banach algebras ${\mathcal{A}}$ and ${\mathcal{B}}(\ell^{2}(X))$ respectively. As a consequence, the holomorphic (Riesz) functional calculi (see e.g. [14, Chap. VII]) of ${\mathcal{A}}$ and ${\mathcal{B}}(\ell^{2}(X))$ coincide. More precisely, if $A\in{\mathcal{A}}$, $U\subseteq\mathbb{C}$ is an open neighborhood of (a connected component of) $\sigma_{{\mathcal{A}}}(A)=\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)$, $\Gamma\subseteq U$ is a contour of $\sigma_{{\mathcal{A}}}(A)=\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)$, and if $z\mapsto f(z)$ is holomorphic on $U$, then
@@
Strictly speaking, the authors in [16] considered algebras satisfying properties (A1--A3) without including property (A0) in their definition. In the language of [16, Definition 2], property (A0) can be rephrased to saying that the constant weight $1$ is ${\mathcal{A}}$-admissible with critical index $p_{0}=1$. Here we assume for simplicity, that a spectral algebra automatically satisfies (A0) as well, since many of the classical examples of spectral algebras do so anyway, see below.
@@
[16, Theorem 3.6] Let $\varphi$ be a frame in $\mathcal{H}$. If $\varphi\sim_{{\mathcal{A}}}\varphi$ then also $\widetilde{\varphi}\sim_{{\mathcal{A}}}\widetilde{\varphi}$ and $\varphi\sim_{{\mathcal{A}}}\widetilde{\varphi}$.
@@
[16, Definition 5] Let $\mathcal{A}$ be a spectral algebra, $\varphi$ be a frame for $\mathcal{H}$, and $\varphi^{d}$ be some dual frame of $\varphi$ such that $\varphi\sim_{{\mathcal{A}}}\varphi^{d}$. Let
@@
[16, Theorem 3.8], [28, Corollary 3.27] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi=(\varphi_{k})_{k\in X}$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then $\mathcal{H}^{\infty}(\varphi)$ is a Banach space with respect to the norm $\|C_{\widetilde{\varphi}}\cdot\|_{\ell^{\infty}(X)}$, and $\|C_{\varphi}\cdot\|_{\ell^{\infty}(X)}$ is an equivalent norm.
@@
[16, Proposition 2.4],[28, Proposition 3.20] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi=(\varphi_{k})_{k\in X}$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then for each $1\leq p<\infty$, the synthesis operator
@@
[16, Theorem 3.8],[28, Corollary 3.27] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi=(\varphi_{k})_{k\in X}$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then for each $1\leq p\leq\infty$, the frame operator $S_{\varphi}=D_{\varphi}C_{\varphi}:\mathcal{H}^{p}(\varphi)\to\mathcal{H}^{p}(\varphi)$ is bounded and invertible.
@@
[6, Proposition 2] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi=(\varphi_{k})_{k\in X}$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then
@@ 3.1. Frame-related Operators
The next two lemmata generalize [20, Proposition 8], where the case of the Jaffard algebra ${\mathcal{A}}=\mathcal{J}_{s}$ is treated. They are also similar to [6, Lemma 6 and 7], except for the frame assumption on $\psi$. In fact, our conclusion is also weaker due to the missing frame assumption, i.e., we cannot conclude here that $C_{\psi}$ has closed range, as explicitly demonstrated in Example 28 (iv).
@@
The next two lemmata are mild generalizations of parts of [6, Lemma 7].
@@ 4. Proof of Theorem 1
which means that $(\omega_{k})_{k\in X}$ is an R-dual sequence of $\big{(}S_{\varphi}^{1/2}\psi_{k}\big{)}_{k\in X}$ in the sense of [12]. By [12, Theorem 2] we know that $(\omega_{k})_{k\in X}$ is a Riesz sequence if and only if $\big{(}S_{\varphi}^{1/2}\psi_{k}\big{)}_{k\in X}$ is a frame, which in turn is equivalent to $(\psi_{k})_{k\in X}$ being a frame, since $S_{\varphi}^{1/2}\in{\mathcal{B}}(\mathcal{H})$ is invertible.
@@ 5. A Comparison to ”Gabor Frames Without Inequalities”
As already mentioned in the introduction, the main insight of [22] (explicitly stated as a corollary therein) was that, for a window function $g$ in the modulation space $M^{1}({\mathbb{R}}^{d})$, the Gabor system $\mathcal{G}={\mathcal{G}}(g,\Lambda)$ over a full-rank lattice $\Lambda\subset\mathbb{R}^{2d}$ being a frame for $L^{2}(\mathbb{R}^{d})$ is equivalent to the injectivity of the analysis operator $C_{{\mathcal{G}}}:M^{\infty}({\mathbb{R}}^{d})\to\ell^{\infty}(\Lambda)$. In contrast to Theorem 1, no closed range condition is listed in this or any other of the conditions from [22, Theorem 3.1]. This is due to the fact that they are automatically fulfilled in this setting, as the next example illustrates.
@@
Indeed, (2$\Rightarrow$(3) is trivial, and (3$\Rightarrow$(1) was proven in [22]. To show (1$\Rightarrow$(2), we note that due to $g\in M^{1}(\mathbb{R}^{d})$, $S_{{\mathcal{G}}}=D_{{\mathcal{G}}}C_{{\mathcal{G}}}$ is bounded and invertible on $M^{p}(\mathbb{R}^{d})$ for all $1\leq p\leq\infty$ [22, Theorem 3.1]. From the latter, we may deduce the desired closed range condition analogously as in the proof of Theorem 1.
In a similar manner, other closed range conditions from [22, Theorem 3.1] can be derived.
At the same time, however, all closed range conditions appearing in Theorem 1 are in fact necessary. To demonstrate this, we ”naively” transfer the items (iii)-(vi), (ix), (x) and (xiii) from [22, Theorem 3.1] to our setting, obtaining the following list of conditions:
@@ 6. Shift-Invariant Spaces
$V^{p}(\varphi)=\left\{f\in L^{p}({\mathbb{R}}^{d}):\ f=\sum_{k\in{\mathbb{Z}}^{d}}c_{k}T_{k}\varphi,\ c\in\ell^{p}({\mathbb{Z}}^{d})\right\}\qquad(1\leq p\leq\infty).$
@@
Then the assumptions (i)--(iii) ensure that each $V^{p}(\varphi)$ is a closed subspace of $L^{p}({\mathbb{R}}^{d})$, and that $V^{2}(\varphi)$ is a reproducing kernel Hilbert space with kernel $K_{x}^{\varphi}$, i.e.
@@
i.e., the Riesz basis $(T_{k}\varphi)_{k\in{\mathbb{Z}}^{d}}$ is intrinsically ${\mathcal{J}}_{s}(\mathbb{Z}^{d})$-localized, where ${\mathcal{J}}_{s}({\mathbb{Z}}^{d})$ denotes the Jaffard algebra from Example 4 (1). Furthermore, from [20] we also know that the associated co-orbit spaces $\mathcal{H}^{p}((T_{k}\varphi)_{k\in\mathbb{Z}^{d}})$ coincide with the shift-invariant spaces $V^{p}(\varphi)$ and
(16) $|\langle K_{x_{k}}^{\varphi},T_{l}\varphi\rangle|=|\varphi(x_{k}-l)|\lesssim(1+|x_{k}-l|)^{-s},\qquad k,l\in\mathbb{Z}^{d}.$
Let $X=(x_{k})_{k\in{\mathbb{Z}}^{d}}\subset{\mathbb{R}}^{d}$ be a (discrete) sampling set. If $\psi=(K_{x_{k}}^{\varphi})_{k\in{\mathbb{Z}}^{d}}$ forms a frame for $V^{2}(\varphi)$ with frame bounds $0<A\leq B$, then $X$ is called a stable set of sampling, since, due to (15), the frame inequality (1) takes the form
$A\|f\|^{2}\leq\sum_{k\in{\mathbb{Z}}^{d}}|f(x_{k})|^{2}\leq B\|f\|^{2},\qquad f\in V^{2}(\varphi)$
in this case.
@@
(17) $|x_{k}-k|\leq C,\qquad\text{for some }C>0\text{ and every }k\in{\mathbb{Z}}.$
@@
$|\langle K_{x_{k}}^{\varphi},T_{l}\varphi\rangle|\lesssim(1+|k-x_{k}|)^{s}(1+|k-l|)^{-s}\lesssim(1+|k-l|)^{-s},\qquad k,l\in\mathbb{Z}^{d},$
@@
which means that $\psi$ and $(T_{k}\varphi)_{k\in\mathbb{Z}^{d}}$ are mutually ${\mathcal{J}}_{s}(\mathbb{Z}^{d})$-localized. Thus all assumptions of Theorem 1 are met in this setting and all of the conditions (1)--(10) provide necessary and sufficient conditions for such a set $X$ being a stable set of sampling.
@@
Finally, we compute the R-dual $\omega$ of $\psi$. Since the frame operator $S_{\varphi}$ commutes with translations by $k\in{\mathbb{Z}}^{d}$, so does $S_{\varphi}^{-1/2}$, and we see that $S_{\varphi}^{-1/2}T_{k}\varphi=T_{k}S_{\varphi}^{-1/2}\varphi=T_{k}\gamma$ for $S_{\varphi}^{-1/2}\varphi=\gamma$ and $k\in{\mathbb{Z}}^{d}$. Consequently,
@@
(18) $\omega_{k}=\sum_{l\in{\mathbb{Z}}^{d}}\langle\psi_{l},T_{k}\varphi\rangle S_{\varphi}^{-{1}/{2}}T_{l}\varphi=\sum_{l\in X}\overline{\varphi(x_{l}-k)}T_{l}\gamma,\qquad k\in{\mathbb{Z}}^{d}.$
@@
Moreover, in case $X=(x_{k})_{k\in{\mathbb{Z}}^{d}}={\mathbb{Z}}^{d}$, the R-dual $\omega$ of $\psi$ is again a family of shifts of a single generator $\eta$ as

### v2

@@ 1. Introduction
In many settings, however, frame theoretic results are ultimately not the consequence of an underlying group structure, but follow from localization properties of the frame, see, e.g., kernel theorems (compare [6, 14] with [10]), invertibility results of Gabor frame operators on modulation spaces [19, 23, 25], or a partial generalization [25, Theorem 5.1] of [21, Theorem 3.1]. A frame $(\psi_{k})_{k\in X}$ is said to be localized, if its Gram matrix is contained in a spectral matrix algebra (see [15] or Definition 2). To a localized frame $\psi$ one can then associate a family of Banach spaces, the so-called co-orbit spaces ${\mathcal{H}}^{p}(\psi)\ (1\leq p\leq\infty)$ which are defined in terms of summability conditions on the analysis coefficients $(\langle f,\psi_{k}\rangle)_{k\in X}$, see Section 2.3 for further details.
Returning to Gabor analysis, we note that the property $g\in M^{1}(\mathbb{R}^{d})$ of the window function can be seen as a localization assumption. Indeed, in the case $\Lambda=\mathbb{Z}^{d}$, we note that ${\mathcal{G}}(g,\mathbb{Z}^{d})$ is localized with respect to the unweighted Baskakov-Gohberg-Sjöstrand algebra $\mathcal{C}$ (see Example 4 (3)), and in the case of more irregular sets $\Lambda$, similar (more general) properties have been shown [25]. Furthermore, if ${\mathcal{G}}$ is a frame, the modulation spaces $M^{p}({\mathbb{R}}^{d})$ coincide with the co-orbit spaces $\mathcal{H}^{p}({\mathcal{G}})$ ($1\leq p\leq\infty$) generated from ${\mathcal{G}}$ in this case [25, Theorem 3.2]. Naturally, these observations lead to the question, whether abstract localized frames can be characterized in a similar fashion as in [21], namely via properties of their associated frame-related operators on their respective co-orbit spaces.
@@
Second, to our knowledge, there is currently no canonical generalization of the Ron-Shen duality principle to abstract frames in the literature. The Ron-Shen duality principle connects any Gabor frame on a lattice with its (dual) Gabor Riesz sequence generated by the same window on the adjoint lattice. In Theorem 1 we consider the family $\omega$, which, in the language of [33], is a so-called R-dual of type III of $\psi$, and serves as a substitute of the aforementioned Ron-Shen dual. In general, however, an R-dual associated with a Gabor frame on a lattice does not coincide with its Ron-Shen dual, see [11, 33] for a discussion of this matter. In order to guarantee that the R-dual $\omega$ is a Riesz sequence if and only if $\psi$ is a frame, we additionally have to assume that the reference system $\varphi$ is in fact a Riesz basis and not only a frame. Thus, our localization assumption $\psi\sim_{\mathcal{A}}\varphi$ corresponds to the original localization concept from [19] in the case ${\mathcal{A}}={\mathcal{J}}_{s}$ (see Example 4 (1)), which was followed by several other publications on this topic [1, 3, 4, 5, 15, 20]. At the same time, however, it might be very difficult, if not even impossible for a Gabor frame to be localized with respect to a localized Riesz basis (see also [19, Section 5] for a discussion of this issue).
Finally, several closed range conditions appear in our setting but not in [21, Theorem 3.1]. We show in Example 28 and Remark 29, that each of the closed range conditions appearing in Theorem 1 is in fact necessary in our setting, since otherwise, the statement is simply not true. For Gabor frames, however, these conditions are automatically satisfied, see Example 27.
By the previous observations, Theorem 1 does not directly generalize [21, Theorem 3.1]. However, [21] served as inspiration for this article, and multiple equivalent conditions of [21, Theorem 3.1] can be translated directly into conditions in Theorem 1. We refer to Section 5 for a more detailed comparison.
This article is structured as follows. In Section 2 we fix our notation and summarize some relevant concepts and results from the literature. In Section 3 we prove some auxiliary results, which we utilize in Section 4, where we prove our main result. In Section 5 we compare Theorem 1 to [21, Theorem 3.1], which initiated the research on the content of this article. Finally, Section 6 is devoted to applying our main result in the setting of shift-invariant spaces.
@@ 2.2. The R-dual sequence
converges in $\mathcal{H}$ for each $k\in X$ and the family $\omega=(\omega_{k})_{k\in X}$ is called an R-dual or Riesz-dual sequence for $\varphi$. It was shown in [11, Propositions 10,11] that $\varphi$ is a frame for ${\mathcal{H}}$ if and only if $\omega$ is a Riesz sequence.
@@ 2.3. Localized Frames
A unital Banach *-algebra $\mathcal{A}$ of $\mathbb{C}$-valued matrices $A=[A_{k,l}]_{k,l\in X}$ is called a solid spectral matrix algebra [15, Section 2.3], or simply spectral algebra, if
@@
Condition (A2) is equivalent to spectral invariance of ${\mathcal{A}}$ in ${\mathcal{B}}(\ell^{2}(X))$, that is, $\sigma_{{\mathcal{A}}}(A)=\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)$ for all $A\in{\mathcal{A}}$ [22, Lemma 2.4], where $\sigma_{{\mathcal{A}}}(A)$ and $\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)$ denote the spectrum of $A$ in the Banach algebras ${\mathcal{A}}$ and ${\mathcal{B}}(\ell^{2}(X))$ respectively. As a consequence, the holomorphic (Riesz) functional calculi (see e.g. [13, Chap. VII]) of ${\mathcal{A}}$ and ${\mathcal{B}}(\ell^{2}(X))$ coincide. More precisely, if $A\in{\mathcal{A}}$, $U\subseteq\mathbb{C}$ is an open neighborhood of (a connected component of) $\sigma_{{\mathcal{A}}}(A)=\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)$, $\Gamma\subseteq U$ is a contour of $\sigma_{{\mathcal{A}}}(A)=\sigma_{{\mathcal{B}}(\ell^{2}(X))}(A)$, and if $z\mapsto f(z)$ is holomorphic on $U$, then
@@
Strictly speaking, the authors in [15] considered algebras satisfying properties (A1--A3) without including property (A0) in their definition. In the language of [15, Definition 2], property (A0) can be rephrased to saying that the constant weight $1$ is ${\mathcal{A}}$-admissible with critical index $p_{0}=1$. Here we assume for simplicity, that a spectral algebra automatically satisfies (A0) as well, since many of the classical examples of spectral algebras do so anyway, see below.
@@
[15, Theorem 3.6] Let $\varphi$ be a frame in $\mathcal{H}$. If $\varphi\sim_{{\mathcal{A}}}\varphi$ then also $\widetilde{\varphi}\sim_{{\mathcal{A}}}\widetilde{\varphi}$ and $\varphi\sim_{{\mathcal{A}}}\widetilde{\varphi}$.
@@
[15, Definition 5] Let $\mathcal{A}$ be a spectral algebra, $\varphi$ be a frame for $\mathcal{H}$, and $\varphi^{d}$ be some dual frame of $\varphi$ such that $\varphi\sim_{{\mathcal{A}}}\varphi^{d}$. Let
@@
[15, Theorem 3.8], [28, Corollary 3.27] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi=(\varphi_{k})_{k\in X}$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then $\mathcal{H}^{\infty}(\varphi)$ is a Banach space with respect to the norm $\|C_{\widetilde{\varphi}}\cdot\|_{\ell^{\infty}(X)}$, and $\|C_{\varphi}\cdot\|_{\ell^{\infty}(X)}$ is an equivalent norm.
@@
[15, Proposition 2.4],[28, Proposition 3.20] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi=(\varphi_{k})_{k\in X}$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then for each $1\leq p<\infty$, the synthesis operator
@@
[15, Theorem 3.8],[28, Corollary 3.27] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi=(\varphi_{k})_{k\in X}$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then for each $1\leq p\leq\infty$, the frame operator $S_{\varphi}=D_{\varphi}C_{\varphi}:\mathcal{H}^{p}(\varphi)\to\mathcal{H}^{p}(\varphi)$ is bounded and invertible.
@@
[5, Proposition 2] Let ${\mathcal{A}}$ be a spectral algebra, and $\varphi=(\varphi_{k})_{k\in X}$ be an ${\mathcal{A}}$-localized frame for $\mathcal{H}$ with canonical dual frame $\widetilde{\varphi}$. Then
@@ 3.1. Frame-related Operators
The next two lemmata generalize [19, Proposition 8], where the case of the Jaffard algebra ${\mathcal{A}}=\mathcal{J}_{s}$ is treated. They are also similar to [5, Lemma 6 and 7], except for the frame assumption on $\psi$. In fact, our conclusion is also weaker due to the missing frame assumption, i.e., we cannot conclude here that $C_{\psi}$ has closed range, as explicitly demonstrated in Example 28 (iv).
@@
The next two lemmata are mild generalizations of parts of [5, Lemma 7].
@@ 4. Proof of Theorem 1
which means that $(\omega_{k})_{k\in X}$ is an R-dual sequence of $\big{(}S_{\varphi}^{1/2}\psi_{k}\big{)}_{k\in X}$ in the sense of [11]. By [11, Theorem 2] we know that $(\omega_{k})_{k\in X}$ is a Riesz sequence if and only if $\big{(}S_{\varphi}^{1/2}\psi_{k}\big{)}_{k\in X}$ is a frame, which in turn is equivalent to $(\psi_{k})_{k\in X}$ being a frame, since $S_{\varphi}^{1/2}\in{\mathcal{B}}(\mathcal{H})$ is invertible.
@@ 5. A Comparison to ”Gabor Frames Without Inequalities”
As already mentioned in the introduction, the main insight of [21] (explicitly stated as a corollary therein) was that, for a window function $g$ in the modulation space $M^{1}({\mathbb{R}}^{d})$, the Gabor system $\mathcal{G}={\mathcal{G}}(g,\Lambda)$ over a full-rank lattice $\Lambda\subset\mathbb{R}^{2d}$ being a frame for $L^{2}(\mathbb{R}^{d})$ is equivalent to the injectivity of the analysis operator $C_{{\mathcal{G}}}:M^{\infty}({\mathbb{R}}^{d})\to\ell^{\infty}(\Lambda)$. In contrast to Theorem 1, no closed range condition is listed in this or any other of the conditions from [21, Theorem 3.1]. This is due to the fact that they are automatically fulfilled in this setting, as the next example illustrates.
@@
Indeed, (2$\Rightarrow$(3) is trivial, and (3$\Rightarrow$(1) was proven in [21]. To show (1$\Rightarrow$(2), we note that due to $g\in M^{1}(\mathbb{R}^{d})$, $S_{{\mathcal{G}}}=D_{{\mathcal{G}}}C_{{\mathcal{G}}}$ is bounded and invertible on $M^{p}(\mathbb{R}^{d})$ for all $1\leq p\leq\infty$ [21, Theorem 3.1]. From the latter, we may deduce the desired closed range condition analogously as in the proof of Theorem 1.
In a similar manner, other closed range conditions from [21, Theorem 3.1] can be derived.
At the same time, however, all closed range conditions appearing in Theorem 1 are in fact necessary. To demonstrate this, we ”naively” transfer the items (iii)-(vi), (ix), (x) and (xiii) from [21, Theorem 3.1] to our setting, obtaining the following list of conditions:
@@ 6. Shift-Invariant Spaces
$V^{p}(\varphi)=\left\{\sum_{k\in{\mathbb{Z}}^{d}}c_{k}T_{k}\varphi:\ c\in\ell^{p}({\mathbb{Z}}^{d})\right\}\qquad(1\leq p\leq\infty).$
@@
Then the assumptions (i)--(iii) ensure that $V^{p}(\varphi)$ is a closed subspace of $L^{p}({\mathbb{R}}^{d})$ ($1\leq p\leq\infty$), and that $V^{2}(\varphi)$ is a reproducing kernel Hilbert space with kernel $K_{x}^{\varphi}$, i.e.
@@
i.e., the Riesz basis $(T_{k}\varphi)_{k\in{\mathbb{Z}}^{d}}$ is intrinsically ${\mathcal{J}}_{s}(\mathbb{Z}^{d})$-localized, where ${\mathcal{J}}_{s}({\mathbb{Z}}^{d})$ denotes the Jaffard algebra from Example 4 (1). Furthermore, from [19] we also know that the associated co-orbit spaces $\mathcal{H}^{p}\big{(}(T_{k}\varphi)_{k\in\mathbb{Z}^{d}}\big{)}$ coincide with the shift-invariant spaces $V^{p}(\varphi)$ ($1\leq p\leq\infty$) and
(16) $|\langle K_{x}^{\varphi},T_{l}\varphi\rangle|=|\varphi(x-l)|\lesssim(1+|x-l|)^{-s},\qquad x\in\mathbb{R}^{d},l\in\mathbb{Z}^{d}.$
Let $X=(x_{k})_{k\in{\mathbb{Z}}^{d}}\subset{\mathbb{R}}^{d}$ be discrete. If $\psi=(\psi_{k})_{k\in{\mathbb{Z}}^{d}}:=(K_{x_{k}}^{\varphi})_{k\in{\mathbb{Z}}^{d}}$ forms a frame for $V^{2}(\varphi)$ with frame bounds $0<A\leq B$, then $X$ is called a stable set of sampling for $V^{2}(\varphi)$, since, due to (15), the frame inequality (1) takes the form
$A\|f\|_{L^{2}(\mathbb{R}^{d})}^{2}\leq\sum_{k\in{\mathbb{Z}}^{d}}|f(x_{k})|^{2}\leq B\|f\|_{L^{2}(\mathbb{R}^{d})}^{2},\qquad f\in V^{2}(\varphi),$
in this case. Similarly, $X$ is called a stable set of sampling for $V^{p}(\varphi)$ ($1\leq p\leq\infty$), if there exist constants $A_{p},B_{p}>0$ such that
$A_{p}\|f\|_{L^{p}(\mathbb{R}^{d})}\leq\|(f(x_{k}))_{k\in{\mathbb{Z}}^{d}}\|_{\ell^{p}({\mathbb{Z}}^{d})}\leq B_{p}\|f\|_{L^{p}(\mathbb{R}^{d})},\qquad f\in V^{p}(\varphi).$
@@
(17) $|x_{k}-k|\leq C,\qquad\text{for some }C>0\text{ and every }k\in{\mathbb{Z}}^{d}.$
@@
$|\langle\psi_{k},T_{l}\varphi\rangle|\lesssim(1+|k-x_{k}|)^{s}(1+|k-l|)^{-s}\lesssim(1+|k-l|)^{-s},\qquad k,l\in\mathbb{Z}^{d},$
@@
which means that $\psi$ and $(T_{k}\varphi)_{k\in\mathbb{Z}^{d}}$ are mutually ${\mathcal{J}}_{s}(\mathbb{Z}^{d})$-localized. Thus all assumptions of Theorem 1 are met in this setting and all of the conditions (1)--(10) provide necessary and sufficient conditions for such a set $X$ being a stable set of sampling. Instead of repeating all these conditions, we will only reformulate conditions (1), (4) and (8)--(10), since they seem more insightful to us in this case than the remaining ones.
@@
Before doing so, we compute the R-dual $\omega$ of $\psi$. Since the frame operator $S_{\varphi}$ commutes with translations by $k\in{\mathbb{Z}}^{d}$, so does $S_{\varphi}^{-1/2}$, and we see that $S_{\varphi}^{-1/2}T_{k}\varphi=T_{k}S_{\varphi}^{-1/2}\varphi=T_{k}\gamma$ for $\gamma:=S_{\varphi}^{-1/2}\varphi$ and $k\in{\mathbb{Z}}^{d}$. Consequently,
@@
(18) $\omega_{k}=\sum_{l\in{\mathbb{Z}}^{d}}\langle\psi_{l},T_{k}\varphi\rangle S_{\varphi}^{-{1}/{2}}T_{l}\varphi=\sum_{l\in{\mathbb{Z}}^{d}}\overline{\varphi(x_{l}-k)}T_{l}\gamma,\qquad k\in{\mathbb{Z}}^{d}.$
@@
Note, that in case $X=(x_{k})_{k\in{\mathbb{Z}}^{d}}={\mathbb{Z}}^{d}$, the R-dual $\omega$ of $\psi$ is again a family of shifts of a single generator $\eta$ as
$\omega_{k}=T_{k}\sum_{l\in{\mathbb{Z}}^{d}}\overline{\varphi(l-k)}T_{l-k}\gamma=T_{k}\sum_{l\in{\mathbb{Z}}^{d}}\overline{\varphi(l)}T_{l}\gamma=T_{k}\eta,\qquad k\in{\mathbb{Z}}^{d}.$
Computing the entries of the associated Gram matrix $G_{\omega}$, we see that